# A Survey of Generative AI for Intelligent Transportation Systems

### v1

@@ 3.1. Traffic Data Imputation
The task of traffic data imputation involves learning a function, denoted as $f(\cdot)$, which takes as input the corrupted data or missing data $X$, in order to reconstruct the complete data $\hat{X}$. Formally, the reconstructed data $\hat{X}$ can be expressed as $\hat{X}=f(X,\Psi)$, where $\Psi$ denotes other important factors like road network graph.
@@
Low-dimensional traffic data. The low-dimensional traffic data like speed, flow and travel time is crucial for various traffic applications such as route planning and traffic prediction. Missing it is a common issue due to sensor device damage or transmission disruptions. This has led researchers to employ advanced generative AI techniques for imputing missing data using available measurements.
VAE. In previous research (Boquet et al., 2019), a VAE was employed to facilitate the imputation of missing traffic data in an online unsupervised manner by drawing from the learned data distribution. Chen et al. (Chen et al., 2021d) proposed a spatio-temporal VAE method for imputing missing data in traffic raster datasets. This method leverages 3D gated convolution and a multi-attention network to effectively capture spatio-temporal dependencies under the VAE architecture.
GAN. Recently, more researchers have adopted GAN-based methods to address this task (Chen et al., 2019; Han et al., 2020b; Kazemi and Meidani, 2021). For example, Chen et al. (Chen et al., 2019) proposed using real or corrupted data as latent codes and introducing representation loss within a GAN model, significantly improving the performance of traffic data imputation. IGANI (Kazemi and Meidani, 2021) introduced a novel iterative GAN imputation architecture, which iterated over imputed data while preserving the reversibility of the generative imputer.
An important characteristic of traffic data lies in its spatiotemporal correlation. Thus, numerous works proposed to incorporate the spatio-temporal modeling in GAN-based architectures (Huang et al., 2021; Yang et al., 2021; Zhang et al., 2021f, d, d; Yang et al., 2022; Huang and Chen, 2022; Shen et al., 2022; Li et al., 2023c, a; Peng et al., 2023). For example, The authors in (Li et al., 2018b) introduced a 3D convolutional GAN for imputing missing traffic data, utilizing a fractional-strided 3D convolutional neural network (CNN) for the generator and a 3D CNN for the discriminator to capture spatio-temporal features efficiently. STGAN (Yuan et al., 2022) incorporated generation and center losses to minimize reconstruction errors and maintain local spatio-temporal distributions, while the discriminator used a CNN classifier to assess global spatio-temporal distribution compliance, and the generator employed skip connections and dilated convolutions to capture spatio-temporal correlations in traffic data. GAE-GAN-LSTM (Xu et al., 2022) leveraged an enhanced graph autoencoder to extract spatio-temporal features, and employed a GAN for generating complete spatio-temporal features from the missing data, with a generator built on long short-term memory and a discriminator using a fully connected neural network.
@@
Diffusion model. Recently, diffusion models have attracted more researchers’ attentions in traffic data imputation (Liu et al., 2023b; Zhang et al., 2023c; Yun et al., 2023). For instance, Liu et al. (Liu et al., 2023b) introduced an enhanced prior modeling conditional diffusion framework for spatio-temporal imputation, which effectively handles various missing patterns and scenarios, including high missing rates and sensor failures, by leveraging conditional feature extraction and noise estimation modules. SaSDim (Zhang et al., 2023c) introduced a self-adaptive noise scaling mechanism based on diffusion model. This model incorporated a new loss function to scale the noise to a consistent intensity level, and a spatial-temporal global convolution module to capture complex spatial-temporal patterns.
@@
CNN, self-attention mechanism (Zhang et al., 2021f)
@@
High-dimensional traffic data. High-dimensional traffic data, such as traffic videos and 3D LiDAR vision, plays a critical role, particularly in the context of autonomous driving, where complete dataset would greatly enhance situational awareness and decision-making capabilities. However, the loss of high-dimensional traffic data remains a prevalent issue in real-world scenarios. To combat this challenge, Wu et al. (Wu et al., 2022) introduced a novel complementary model based on GANs for addressing missing frames in traffic videos. The model utilized a feature pyramid network to extract multi-scale feature maps from input video frames, enhancing the integration of semantic information across frames by fusing features from different scales. The inclusion of local block discriminators within the discriminator model effectively ensured the accuracy and continuity of the generated frames. TDC-GAN (Wu et al., 2021) utilized a multi-scale semantic information extraction model based on the feature pyramid network and a discriminator model with global and local branches to ensure spatio-temporal consistency in consecutive frames. For incomplete point clouds in LiDAR applications, Tu et al. (Tu et al., 2021) introduced a GAN-based point cloud repair network specifically applied within the autonomous driving context, enhancing overall safety in autonomous driving systems.
@@ 3.2. Traffic Estimation
Several works designed GAN-based approaches to estimate traffic condition (Liang et al., 2018; Yu and Gu, 2019; Zhang et al., 2021c; Mo et al., 2022; Tian et al., 2022; Zhang et al., 2022c; Xu et al., 2023c). GAA (Liang et al., 2018) incorporated traffic-flow theory into a deep GAN model, while GCGA (Yu and Gu, 2019) combined the graph modeling capabilities of GCN and the generative capabilities of GAN . Curb-GAN (Zhang et al., 2020b) enhanced the conditional GAN structure by modeling various travel demands. GE-GAN (Xu et al., 2020b) learns road network features through graph embedding and applies them to a GAN network to estimate traffic conditions. PA-GAN (Tian et al., 2022) incorporated Bayesian inference to adaptively adjust parameters for various traffic patterns based on contextual features, while also utilizing GAN to learn these traffic patterns from sparsely sampled data. Zhang et al. (Zhang et al., 2021c) models the intricate correlations between road traffic and urban conditions in their proposed conditional GAN-based traffic estimation model. Unlike them, the authors in (Yuan et al., 2023) introduced parameterization of noise factors in a diffusion-based model, transforming the traffic matrix estimation problem into a gradient-descent optimization task.
@@ 3.3. Traffic Data Mining
Trajectories. Trajectories refer to a sequence of timestamped positions that describe the movement or path of an object over time, which provides valuable insights for analyzing various aspects such as urban transportation modes and human decision-making preferences. To identify the transportation modes, the authors in (Zhang et al., 2022a) introduced a semi-supervised method based on the Dirichlet VAE, which fuses geographic information with motion features from GPS trajectories. Zhang et al. (Zhang et al., 2020d) proposed a trajectory GAIL framework to learn human decision-making behaviors. This framework modeled the human decision processes using Markov decision processes with varying lengths to capture long-term decision dependencies. Further, it inferred the decision-making strategy from the historical dataset of the human agent in an inverse manner.
@@
Traffic images. In autonomous driving scenarios, traffic images captured by cameras offer vital information about the surrounding driving environment, enabling autonomous vehicles to make accurate decisions. To overcome the challenge of limited traffic image diversity in real urban environments, Li et al. (Li et al., 2020b) introduced a GAN-based framework with structural information. This approach enables various urban image transformations while maintaining the integrity of foreground objects and image structural details. Zhao et al. (Zhao et al., 2022a) focused on the quality of an in-vehicle image. They proposed a GAN-based single-image raindrop removal network that combines task-specific visual attention and content perception mechanisms, allowing the network to focus on raindrop regions and their surroundings for effective raindrop removal.
@@
Unlike them, some researchers focus on extracting valuable information from traffic images for tasks such as vehicle detection (Wang et al., 2021b), license plate identification (Lee et al., 2018; Han et al., 2020a; Ibrahim et al., 2022; Boby et al., 2023), traffic density recognition (Gawali and Deore, 2023), driving behaviors (Wijaya et al., 2022) and map enrichment (Zhong et al., 2022). These tasks often face challenges related to low resolution and inadequate lighting conditions. For example, the authors in (Hassan Onim et al., 2023) presented a novel vehicle detection approach that utilizes two different GANs to address low-light traffic scenarios. Boby et al. (Boby et al., 2023) conducted a comparative analysis between a GAN-based super-resolution model and an iterative refinement approach utilizing diffusion-based super-resolution models for license plate detection. In addition, the authors in (Gawali and Deore, 2023) presented a dual-discriminator conditional GAN for recognizing traffic density in both homogeneous and heterogeneous traffic scenarios.
Traffic videos. Compared to traffic images, traffic videos provide a wealth of redundant traffic information for autonomous vehicles, guaranteeing both safety and efficiency in navigation. A pivotal stage in numerous intelligent video processing applications, such as vehicle behavior analysis, is the segmentation of moving objects. RMS-GAN (Patil et al., 2020) is an end-to-end GAN using a recurrent technique, which integrates foreground probability information using residual and weight-sharing techniques to achieve precise segmentation. Krishna et al. (Krishna and Denzler, 2014) progressively combined generative and discriminative models to efficiently classify activities in traffic videos. Cheng et al. (Cheng et al., 2021a) proposed a deep conditional generative model to detect interactions between vehicles and vulnerable road users. This model utilized a conditional VAE-based framework with Gaussian latent variables to capture the behavior of road users and predict interactions in a probabilistic manner.
@@
Vehicle classification GAN (Wang et al., 2021b)
@@
Traffic video Vehicle-road user interaction detection VAE (Cheng et al., 2021a)
@@ 3.4. Traffic Anomaly Detection
Traffic anomalies refer to unusual events, patterns, or behaviors within a transportation system. These anomalies can disrupt the regular traffic flow or pose potential safety risks. Numerous studies have explored the use of generative AI techniques for traffic anomaly detection (Qiu et al., 2019; Lin et al., 2020; Kumar et al., 2020; Xie et al., 2023). Chen et al. (Chen et al., 2021b) presented a multi-modal GAN model to detect and classify the traffic events, leveraging data from multiple modalities to enhance detection accuracy. The authors in (Santhosh et al., 2021) proposed a hybrid architecture that combines CNN and VAE to detect and classify anomalies. Different from the above GAN-based methods, the work by Kang et al. (Kang et al., 2022) employed conditional normalizing flow to detect abnormal temporal snapshots, followed by road segment-level anomaly detection using a kernel density estimator. A-VAE (Aslam and Kolekar, 2023) was designed as an attention-based VAE architecture for anomaly detection in traffic videos. It utilized a combination of 2D CNN and BiLSTM layers integrated with an attention network for effective representation learning.
Cyberattacks would pose significant safety risks in the context of automated driving. To address it, Zhao et al. (Zhao et al., 2022c) presented a GAN-based intrusion system for connected vehicles to evaluate the legitimacy of incoming messages. Ding et al. (Ding et al., 2019) investigated the susceptibility of deep generative models in autonomous driving to Trojan attacks, revealing that manipulating training data and injecting carefully crafted data can instill Trojan behaviors in a deep generative model without altering its original training objective. Li et al. (Li et al., 2022c) identified two potential cyberattack scenarios: individual vehicle-targeted attacks and sensor measurement data injection attacks. To address these threats, they introduced a real-time anomaly detection model using GANs for detection purposes.

### v2

@@ 3.1. Traffic Data Imputation
Traffic data imputation involves learning a function, $f(\cdot)$, that takes corrupted or missing data $X$ as input to reconstruct the complete data $\hat{X}$. This can be expressed as $\hat{X}=f(X,\Psi)$, where $\Psi$ represents additional factors like the road network graph.
@@
Low-dimensional traffic data. Low-dimensional traffic data, such as speed, flow, and travel time, is crucial for various applications like route planning and traffic prediction. However, missing data is common due to sensor failures or transmission issues. To address this, researchers have employed advanced generative AI techniques to impute these missing data.
VAE. Boquet et al.(Boquet et al., 2019) developed an online unsupervised VAE-based method for imputing missing traffic data, enhancing forecasting accuracy while enabling data augmentation, compression, and classification. Building on this, Chen et al.(Chen et al., 2021d) introduced the STVAE model, which incorporates 3D gated convolutions, a multi-attention network, and Sylvester normalizing flows to better capture complex spatio-temporal dependencies.
GAN. Recently, more researchers have adopted GAN-based methods to address this task (Chen et al., 2019; Han et al., 2020b; Kazemi and Meidani, 2021). For example, Chen et al. (Chen et al., 2019) proposed using real or corrupted data as latent codes and introducing representation loss within a GAN model. IGANI (Kazemi and Meidani, 2021) introduced a novel iterative GAN imputation architecture, which iterated over imputed data while preserving the reversibility of the generative imputer. In particular, numerous works proposed to incorporate the spatio-temporal modeling in GAN-based architectures (Huang et al., 2021; Yang et al., 2021; Zhang et al., 2021e, c, c; Yang et al., 2022; Huang and Chen, 2022; Shen et al., 2022; Li et al., 2023c, a; Peng et al., 2023). For example, the authors in (Li et al., 2018) introduced a 3D convolutional GAN that excels in capturing intricate traffic dynamics through fractionally strided and standard 3D CNNs for the generator and discriminator, respectively. STGAN (Yuan et al., 2022) enhanced GANs with generation and center losses, along with skip connections and dilated convolutions, achieving superior imputation while maintaining data consistency. GAE-GAN-LSTM (Xu et al., 2022) integrated a graph autoencoder with GANs and LSTM, offering robust imputation across various loss ratios by effectively modeling complex traffic patterns. These innovations highlight the superior capability of advanced GAN-based methods in capturing and reconstructing the complex spatio-temporal relationships in traffic data, offering more reliable and versatile solutions.
@@
DPM. Recently, diffusion models have attracted more researchers’ attentions in traffic data imputation (Liu et al., 2023b; Zhang et al., 2023c; Yun et al., 2023). For instance, Liu et al. (Liu et al., 2023b) introduced PriSTI, an enhanced conditional diffusion framework that handles diverse missing patterns and high missing rates by integrating conditional feature extraction and advanced noise estimation. This method mitigates error accumulation seen in autoregressive models and incorporates geographic relationships, offering a more context-aware imputation. Similarly, SaSDim (Zhang et al., 2023c) enhanced diffusion models with a self-adaptive noise scaling mechanism and a spatial-temporal global convolution module. Their novel loss function normalizes noise intensity, reducing the negative impact of noise and capturing dynamic dependencies more effectively. These methods illustrate the potential of diffusion models to address spatio-temporal imputation tasks by leveraging their inherent flexibility in handling noise and capturing complex data dependencies.
@@
CNN, self-attention mechanism (Zhang et al., 2021e)
@@
High-dimensional traffic data. High-dimensional traffic data, such as traffic videos and 3D LiDAR vision, plays a crucial role in enhancing situational awareness and decision-making, particularly in autonomous driving. However, the loss of high-dimensional data remains a common challenge in real-world applications. To address this, Wu et al. (Wu et al., 2022, 2021) utilized feature pyramid networks within GAN frameworks to extract multi-scale semantic features from traffic videos. This multi-scale approach enhanced the integration of spatial and temporal information, ensuring continuity and accuracy in the imputed frames. The incorporation of local and dual-branch discriminators further guaranteed spatio-temporal consistency, which was crucial for applications requiring precise motion and context understanding, such as traffic monitoring and autonomous navigation. Tu et al. (Tu et al., 2021) extended GAN applications to 3D LiDAR point clouds with their Point Fractal Network (PF-Net). By repairing incomplete point clouds, PF-Net restores vital geometric and semantic information, thereby improving object detection and collision avoidance in autonomous vehicles. This GAN-based repair not only enhances data integrity but also supports the real-time requirements of autonomous driving systems.
@@ 3.2. Traffic Estimation
Several works designed GAN-based approaches to estimate traffic condition (Liang et al., 2018; Yu and Gu, 2019; Zhang et al., 2021b; Mo et al., 2022; Tian et al., 2022; Zhang et al., 2022c; Xu et al., 2023c). GAA (Liang et al., 2018) merged traffic-flow theory with GANs using LSTM networks to effectively capture spatio-temporal correlations. This integration not only improved estimation accuracy compared to Bayesian methods but also demonstrated the potential of combining theoretical models with deep learning. GCGA (Yu and Gu, 2019) combined GCN with GANs, capturing the spatial dependencies and leading to more accurate traffic condition estimations. Innovations like Curb-GAN (Zhang et al., 2020b) and GE-GAN (Xu et al., 2020b) further refined these models by incorporating dynamic convolutional layers and graph embeddings, enabling more precise and context-aware traffic estimations. PA-GAN (Tian et al., 2022) integrated Bayesian inference to adaptively adjust parameters based on contextual features, effectively handling multi-modal traffic patterns and reducing estimation biases. Additionally, Zhang et al. (Zhang et al., 2021b) models the complex correlations between road traffic and urban conditions in their proposed conditional GAN-based model, further improving estimation accuracy.
Unlike them, the authors in (Yuan et al., 2023) leveraged denoising DPM by parameterizing noise factors, transforming traffic matrix estimation into a gradient-descent optimization task. This method demonstrated superior performance in both traffic matrix synthesis and estimation. Lei et al.(Lei et al., 2024) further extended diffusion models with their Conditional Diffusion Framework with Spatio-Temporal Estimator (CDSTE), which integrates probabilistic diffusion with spatio-temporal networks. CDSTE excels in inferring complete traffic states at sensor-free locations, providing reliable probabilistic estimates that support robust traffic management and control decisions.
@@ 3.3. Traffic Data Mining
Trajectories. Trajectories are sequences of timestamped positions that describe the movement of an object over time, providing valuable insights for analyzing various aspects such as urban transportation modes and human decision-making preferences. To identify the transportation modes, the authors in (Zhang et al., 2022a) introduced a semi-supervised Dirichlet VAE that integrates geographic data with motion features. This model effectively leveraged limited labeled GPS data and accounts for geographic factors, achieving superior classification accuracy on real-world datasets. Similarly, TrajGAIL (Zhang et al., 2020d) used Variable Length Markov Decision Processes and deep neural networks to capture long-term dependencies in human decisions. This model outperformed traditional Markov-based models, demonstrating enhanced accuracy in replicating complex behaviors.
@@
Traffic images. In autonomous driving scenarios, traffic images captured by cameras provide essential information about the surrounding environment, allowing autonomous vehicles to make effective decisions. To overcome the challenge of limited image diversity in real urban environments, Li et al. (Li et al., 2020b) introduced a GAN-based framework that performs diverse urban image style transfers, such as day-to-night transitions, while effectively preserving foreground objects and structural details. This method outperformed baseline models by eliminating distortions and maintaining high-quality transformations, thereby increasing the variability and reliability of training data for autonomous systems. Zhao et al. (Zhao et al., 2022a) addressed image degradation from raindrops on in-vehicle cameras with a GAN-based removal network that integrates visual attention and content perception. This approach not only enhances image quality but also improves traffic object detection accuracy.
@@
Unlike them, some researchers focus on extracting valuable information from traffic images for tasks such as vehicle detection (Wang et al., 2021a), license plate identification (Lee et al., 2018; Han et al., 2020a; Ibrahim et al., 2022; Boby et al., 2023), traffic density recognition (Gawali and Deore, 2023), driving behaviors (Wijaya et al., 2022) and map enrichment (Zhong et al., 2022). These tasks often face challenges related to low resolution and inadequate lighting conditions. For example, the authors in (Hassan Onim et al., 2023) enhanced vehicle detection in low-light conditions by integrating Pix2PixGAN and CycleGAN with a YOLOv4-based algorithm. This method effectively overcomes the limitations of traditional vision techniques in poor lighting, demonstrating substantial performance improvements that enhance autonomous vehicle reliability. In the domain of license plate detection, Boby et al. (Boby et al., 2023) compared a GAN-based super-resolution model with diffusion-based iterative refinement models. This comparison highlighted the importance of selecting the appropriate super-resolution technique based on specific application requirements, with GANs offering significant advantages in measurable performance metrics. In addition, the authors in (Gawali and Deore, 2023) presented a dual-discriminator conditional GAN for recognizing traffic density in both homogeneous and heterogeneous traffic scenarios. The dual-discriminator architecture, optimized using Giza pyramids construction, enhanced the model’s ability to distinguish between varying traffic densities with improved accuracy and reduced loss.
@@
Vehicle classification GAN (Wang et al., 2021a)
@@
Traffic video Vehicle-road user interaction detection VAE (Cheng et al., 2021)
@@
Traffic videos. Compared to traffic images, traffic videos provide a wealth of redundant traffic information for autonomous vehicles, guaranteeing both safety and efficiency in navigation. A pivotal stage in numerous intelligent video processing applications, such as vehicle behavior analysis, is the segmentation of moving objects. RMS-GAN (Patil et al., 2020) is an end-to-end GAN using a recurrent technique, which integrates foreground probability information using residual and weight-sharing techniques to achieve precise segmentation. Krishna et al. (Krishna and Denzler, 2014) progressively combined generative and discriminative models to efficiently classify activities in traffic videos. Cheng et al. (Cheng et al., 2021) proposed a deep conditional generative model to detect interactions between vehicles and vulnerable road users. This model utilized a conditional VAE-based framework with Gaussian latent variables to capture the behavior of road users and predict interactions in a probabilistic manner. DiffBEV (Zou et al., 2024) addressed the issue of noisy Bird’s Eye View (BEV) representations in autonomous driving by using a conditional diffusion model to improve BEV feature quality through progressive refinement. The authors in (Fang et al., 2024a) proposed an abductive framework AdVersa-SD for accident video comprehension. AdVersa-SD employs an object-centric video diffusion approach to learn cause-effect relationships of accidents, thereby enhancing safe driving perception.
@@ 3.4. Traffic Anomaly Detection
Traffic anomalies refer to unusual events, patterns, or behaviors within a transportation system. These anomalies can disrupt the regular traffic flow or pose potential safety risks. Numerous studies have explored the use of generative AI techniques for traffic anomaly detection (Qiu et al., 2019; Lin et al., 2020; Kumar et al., 2020; Xie et al., 2023). Chen et al. (Chen et al., 2021b) presented a multi-modal GAN model to detect and classify the traffic events, leveraging data from multiple modalities to enhance detection accuracy. The authors in (Santhosh et al., 2021) proposed a hybrid CNN-VAE architecture that employs color gradient representations and semi-supervised annotation, achieving a 1-6% increase in classification accuracy and a 30-35% boost in anomaly detection through effective feature extraction and semi-supervised labeling. The work by Kang et al. (Kang et al., 2022) addressed congestion anomaly detection by integrating conditional normalizing flows with kernel density estimators to effectively capture spatial-temporal dependencies in multivariate time series data. Their framework not only demonstrated superior performance, but also ensured robustness across diverse traffic scenarios. A-VAE (Aslam and Kolekar, 2023) integrated 2D CNN and BiLSTM layers with an attention mechanism for enhanced representation learning in traffic video anomaly detection. This approach demonstrated strong real-time processing capabilities, highlighting its practical applicability for live traffic monitoring systems. Li et al. (Li et al., 2024a) presented DiffTAD, a diffusion model-based framework for vehicle trajectory anomaly detection that addresses training instability and mode collapse issues prevalent in GANs and VAEs. By modeling anomaly detection as a noisy-to-normal process and incorporating Transformer-based encoders for temporal and spatial dependencies, DiffTAD achieved superior performance and robustness in diverse traffic environments. Lastly, Yan et al. (Yan et al., 2023) introduced a novel diffusion model–based method for video anomaly detection that utilizes two denoising diffusion implicit modules to predict and refine the motion and appearance features of video frames. The proposed method effectively learned the distribution of normal samples, enabling accurate detection of anomalies.
Cyberattacks would pose significant safety risks in the context of automated driving. To address it, researchers have proposed a variety of solutions leveraging generative models. Zhao et al. (Zhao et al., 2022c) presented GVIDS, a GAN-based intrusion detection system for connected vehicles that distinguishes legitimate CAN messages from malicious ones. Although GVIDS is suitable for dynamic vehicular environments, its reliance on specific CAN message characteristics may limit adaptability to new attack types. Ding et al. (Ding et al., 2019) exposed the vulnerability of deep generative models in autonomous driving to Trojan attacks through data poisoning. Their work underscored the critical risk of maliciously implanted behaviors that activate under specific triggers, highlighting the need for robust data integrity measures. Li et al. (Li et al., 2022c) developed a GAN-based real-time anomaly detection model capable of identifying subtle adversarial manipulations in vehicle and sensor data, demonstrating its potential for dynamic environments. In addition, the authors in (Devika et al., 2024) presented an unsupervised GAN-based framework optimized with LSTM networks to analyze vehicle dynamics, effectively detecting eleven distinct attack types. By integrating LSTM architectures, the framework effectively handled extensive vehicular data, demonstrating robustness in maintaining the integrity of connected and autonomous vehicle networks.
3.5. Discussion
We have summarized the key insights and challenges of generative AI in traffic perception. Generative AI techniques such as VAE, GAN, and DPM have shown advantages in handling missing low- and high-dimensional traffic data, particularly in capturing complex spatiotemporal dependencies, improving traffic prediction and management. However, their ability to handle data noise and dynamic changes varies, requiring scenario-specific selection. Generative AI has also enhanced traffic estimation by integrating traditional traffic theory and deep learning, significantly improving accuracy. In traffic data mining, progress has been made in trajectory analysis, image processing, and video analysis, but challenges related to data diversity and model generalizability exist. In anomaly detection, generative AI effectively identifies both regular and anomalous patterns, but real-time performance and adaptability need improvement. Future research should focus on optimizing spatiotemporal modeling, improving data robustness, and integrating integration with traditional traffic theory to optimize ITS.